OpenStudy (alexandervonhumboldt2):

Prove that 6^18 + 36^20 can be divided by 37. No calculators and dont solve the equation.

OpenStudy (alexandervonhumboldt2):

@ganeshie8 @perl @texaschic101 @ParthKohli

OpenStudy (thomas5267):

\(6^{18}\div36^{20}\) is not a integer...

OpenStudy (alexandervonhumboldt2):

ahh + not /

ganeshie8 (ganeshie8):

\[6^{18} + 36^{20} = 36^9 + 36^{20}\equiv (-1)^9 +(-1)^{20} \equiv -1+1\equiv 0 \pmod{37}\]

OpenStudy (kainui):

Well 37 is 6^2+1 I suppose that would be a good starting place, I don't know.

OpenStudy (alexandervonhumboldt2):

hmmm

ganeshie8 (ganeshie8):

if mods dont make sense, start with kainui's step..

OpenStudy (alexandervonhumboldt2):

(6^18+6^40)/(6^2+1)

OpenStudy (kainui):

ganeshie knows more about these kinds of things than me. In my opinion all numbers are divisible by 37 lol...

OpenStudy (mathmath333):

ganeshie's answer is optimal here

ganeshie8 (ganeshie8):

you may try below if you're familiar with binomial thm \[6^{18}+36^{20} = 36^9 + 36^{20} \\= (37-1)^9 + (37-1)^{20} \\=37M + (-1)^9 + 37N + (-1)^{20} \\= 37X -1 + 1\]

OpenStudy (alexandervonhumboldt2):

what is M?

ganeshie8 (ganeshie8):

good question :) that means you have followed upto 3rd line ?

OpenStudy (alexandervonhumboldt2):

uhhh

OpenStudy (xapproachesinfinity):

hmm interesting! the way you did @ganeshie8

OpenStudy (perl):

"prove that 6^18 + 36^20 is divisible by 37. " This is equivalent to finding an integer k such that 6^18 + 36^20 = 37k

OpenStudy (alexandervonhumboldt2):

so m is instead of k?

OpenStudy (xapproachesinfinity):

tht M is supposed to be all terms that were multiples by 37 when you do binomial theorem or expansion

OpenStudy (alexandervonhumboldt2):

ok so of course 37M divides by 37

OpenStudy (alexandervonhumboldt2):

awesome work @ganeshie8 thx

OpenStudy (alexandervonhumboldt2):

but wait how (37-1)^9=37M-1^9?

OpenStudy (alexandervonhumboldt2):

soo the final would be 36*37*(37^10+e)?

OpenStudy (perl):

let me be more rigorous, make this airtight

OpenStudy (perl):

the mod way works too , that ganeshie suggested

OpenStudy (alexandervonhumboldt2):

hmmmm how 36^9+36^20=36*(1+36^11)? mybe 36^9*(1+36^11)?

OpenStudy (alexandervonhumboldt2):

2sd and 3rf line

OpenStudy (perl):

typo

OpenStudy (alexandervonhumboldt2):

thx

OpenStudy (alexandervonhumboldt2):

tthx

OpenStudy (perl):

6^18 + 36^20 = (6^2)^9 + 36^20 =36^9 + 36^20 =36^9*(1 + 36^11) = 36^9*( 1 + (37-1) ^11 ) So by binomial theorem = 36^9( 1 + 37^11 + terms*37 - 1^11 ) = 36^9( 1 + 37^11 + terms*37 -1) = 36^9( 37^11 + terms*37 ) = 36^9*37*( 37^10 + terms*37) so thats a multiple of 37

OpenStudy (perl):

and to be sure about the binomial theorem part. look at the expansion of a binomial expression (a + b)^n , show that terms are multiple of the first term 'a', except for the last term